OpenStudy (anonymous):

(9x + 2) – 3(x – 4) = 6 help is needed

Nnesha (nnesha):

distribute 2nd parenthesis by -3

OpenStudy (anonymous):

ok thank you @Nnesha

Nnesha (nnesha):

hmm got it ??

Nnesha (nnesha):

do you know how to distribute ??

OpenStudy (anonymous):

ya now i got it thank you

Nnesha (nnesha):

okay :)

OpenStudy (macielynn):

9x+2-3x-4=6 combine like terms 6x+2-4=6 6x+2=6 add 2 to both sides 6x=8 divide by 6 on both sides x=1 1/3

OpenStudy (anonymous):

thank you @macielynn